OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x^2 + 8x − 4.

OpenStudy (anonymous):

your x value would be -b/2a Solve for y with your x value

OpenStudy (anonymous):

Because that makes sense..

OpenStudy (anonymous):

I don't get it ..

OpenStudy (anonymous):

b= coefficient of x= 8 a = coefficient of x^2 = 2

OpenStudy (anonymous):

Make sense now?

OpenStudy (anonymous):

ok stop

OpenStudy (anonymous):

what is a coefficient?

OpenStudy (anonymous):

ok 8x The coefficient is the number that comes before it. Which is 8. \[2x^2\] Your coefficient here would be 2.The number before it. Make sense?

OpenStudy (anonymous):

... not really.

OpenStudy (anonymous):

2x^2 + 8x - 4. I have to find the vertex.

OpenStudy (anonymous):

So, what would the coefficients have to do with anything?

OpenStudy (anonymous):

ok The number before a function or an unknown value x, a, b, n is called a coefficient. If I have \[x^2 + 76x+87y+900n\] My coefficients are 1, from x^2 76 from x 87 from y and 900 from n.

OpenStudy (anonymous):

You're not paying attention are you?

OpenStudy (anonymous):

I'm reading.

OpenStudy (anonymous):

-b/2a b and 2a ARE COEFFICIENTS FROM YOUR EQUATION. The x value of your vector= -b/2a Your y value is f(-b/2a)

OpenStudy (anonymous):

Trying to get it, okay I get that part.

OpenStudy (anonymous):

.. I don't get the vector part & f(-b/2a)

OpenStudy (anonymous):

Is there a formula you could give me for this?

OpenStudy (anonymous):

Your equation is in this form \[ax^2+bx+c\] a is the coefficient of x^2 b is the coefficient of x and c is a constant.

OpenStudy (anonymous):

Ok..

OpenStudy (anonymous):

I gave you the formula in my FIRST comment -b/2a = x value then plug this value that you get into your equation to get the y value. Your vertex is in this form (x,y)

OpenStudy (anonymous):

why -b?

OpenStudy (anonymous):

because that's the formula

OpenStudy (anonymous):

so it'd be -8/2a?

OpenStudy (anonymous):

a = -4?

OpenStudy (anonymous):

a=2 coefficient of x^2 -4 is c

OpenStudy (anonymous):

so your x value would be -8/2*2

OpenStudy (anonymous):

so 8

OpenStudy (anonymous):

\[-8/(2\times2)\]

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

okay so the formula for this would be? I'm going to write it down so I know how to do these.

OpenStudy (anonymous):

yes that is one value of your vertex. The x value. Now plug in -2 into your equation to find the y value and your answer should be in the form (x,y)

OpenStudy (anonymous):

soo.. y = 2(-2)^2 + 8(-2) − 4.

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

-4^2 = 16 y = 16 + -16 - 4

OpenStudy (anonymous):

y = -4

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

now back to the beginning

OpenStudy (anonymous):

we have a problem

OpenStudy (anonymous):

No its -12

OpenStudy (anonymous):

my options to choose from are (−2, 20) (−2, −12) (2, 20) (−2, −28)

OpenStudy (anonymous):

how is it -12

OpenStudy (anonymous):

2(4)+8(-2)-4= 8-16-4= -12

OpenStudy (anonymous):

well. I don't get it.. once again..

OpenStudy (anonymous):

Ok what don't you get?

OpenStudy (anonymous):

The whole process.

OpenStudy (anonymous):

ok do you remember how we got x?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

-b/2a

OpenStudy (anonymous):

-b/2a= x

OpenStudy (anonymous):

where b and a are coefficients. Remember this. \[ax^2 +bx +c\] a is the number before x^2 (coefficient) b is the number before x (coefficient) and c is a constant. The number that's on it's own.